OpenStudy (anonymous):

Find the slope of the line. (-2,3) and (5,3) A.1/7 B.-1/7 C.-7 D.7

OpenStudy (karatechopper):

y1-y2/x1-x2

OpenStudy (karatechopper):

Slope formula^

OpenStudy (anonymous):

to find slope given two points given\[(x_1,y_1) and (x_2,y_2)\] \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] m is the slope

OpenStudy (anonymous):

so using that, what do you get for your slope......

OpenStudy (anonymous):

??

OpenStudy (nubeer):

@jimmyjoew ?

OpenStudy (anonymous):

3-3=0 -2-5=-7

OpenStudy (anonymous):

so is c

OpenStudy (anonymous):

if that is the case, then \[m=\frac{ 0 }{ 7 } = 0\ i think there is something wrong with the points you gave above..(a typo maybe)

OpenStudy (anonymous):

are you there??

OpenStudy (anonymous):

i put everything that i have

OpenStudy (anonymous):

what u think i should put

OpenStudy (anonymous):

okay....so if that is what you were given, then your slope is zero...as we worked it out...

OpenStudy (anonymous):

ok thankx

OpenStudy (anonymous):

yw....(: